A body is acted upon by a force which is proportional of the distance covered. If distance covered be denoted by x,then work done by the force will be proportional to?

Grade:12

1 Answers

Ritu Bisht
7 years ago
since force\alpha f\alphaxis proportional to x
i.e. maf\alphax\alpha=kx where k is any constant
      mvdv/dx=kx
       vdv=k’xdx where k’=k/m
        on integrating we have 
         v2/2= k’x2/2
          v=k’x
         work done = change in kinetic energy
                         i.e. proportional to x2